OpenStudy (lilai3):

Geography: At an elevation of -52 feet, Lake Eyre is the lowest point in Australia. HOw much lower than Lake Eyre is the Valdes Peninsula in South America, which has an elevation of -131 ft?

OpenStudy (anonymous):

-131- (-52)

OpenStudy (lilai3):

is that suppose to be written as an integer or not?

OpenStudy (anonymous):

you calc it :P

OpenStudy (anonymous):

-131 - (-52)= -131+52=-79

OpenStudy (lilai3):

okay thanks

OpenStudy (anonymous):

np